Crew Sleep Cycle Adjustments: Crew wake/sleep cycle today shifted 3.5 hrs to the right for the next days: wakeup - 5:30am, sleep - 9:00pm EDT.
Upon wake-up, Sergey Volkov terminated his fifth MBI-12 SONOKARD experiment session for the long-term Russian sleep study, by taking the recording device from his SONOKARD sports shirt pocket and later copying the measurements to the RSE-MED laptop for subsequent downlink to the ground. (SONOKARD objectives are stated to (1) study the feasibility of obtaining the maximum of data through computer processing of records obtained overnight, (2) systematically record the crewmember's physiological functions during sleep, (3) study the feasibility of obtaining real-time crew health data. Investigators believe that contactless acquisition of cardiorespiratory data over the night period could serve as a basis for developing efficient criteria for evaluating and predicting adaptive capability of human body in long-duration space flight.)
The crew performed the regular weekly three-hour task of thorough station cleaning. ("Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the Service Module (SM) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE's sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.)
As part of the house cleaning, the crew also conducted regular maintenance inspection & cleaning on fan grilles in the FGB (TsV2), SM (VPkhO & VPrK), DC1 (V1, V2, V3) screens and PF1/PF2 dust filters, as well as in the COL (Columbus Orbital Laboratory).
Gregory Chamitoff filled out the regular FFQ (Food Frequency Questionnaire), his third, on the MEC (Medical Equipment Computer). (On the FFQs, NASA astronauts keep a personalized log of their nutritional intake over time on special MEC software. Recorded are the amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins. The FFQ is performed once a week to estimate nutrient intake from the previous week and to give recommendations to ground specialists that help maintain optimal crew health. Weekly estimation has been verified to be reliable enough that nutrients do not need to be tracked daily.)
At ~10:10am EDT, the crewmembers held their regular WPC (Weekly Planning Conference) with the ground, discussing next week's "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.
The three crewmembers conducted their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR), RED resistive exercise device (CDR, FE-1, FE-2), and VELO bike with bungee cord load trainer (CDR, FE-1).
Afterwards, Chamitoff transferred the crew's exercise data file to the MEC for downlink, as well as the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).
The Russian crewmembers had their weekly PFCs (Private Family Conferences) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop), Sergey at ~11:00am, Oleg at ~11:45am.
Working off his suggested 'time permitting' task list, CDR Volkov performed another session of the Russian GFI-8 "Uragan" (hurricane) earth-imaging program using the HVR-Z1J SONY digital cameral. (Uploaded target today was Darwin Island (-92.0E; 1.7S), one of the islands in the Galapagos Archipelago.)
For FE-1 Kononenko, the Russian voluntary task list suggested another ECON KPT-3 session for today, making observations and taking aerial photography of environmental conditions for Russia's Environmental Safety Agency (ECON) using the Nikon D2X digital camera with SIGMA 300-800mm telephoto lens.
CDRA Activity: The US Carbon Dioxide Removal Assembly was activated by ground commanding at ~9:00am for two half-cycles, until ppCO2 had dropped below ~4.5 mmHg, then turned off again. Crew support consisted in deactivation of the ARS (Atmosphere Revitalization Systems) rack.
Airlock (A/L) CCAA Activity: Also by ground commanding, the THC CCAA (Temperature & Humidity Control/Common Cabin Air Assembly) air conditioner in the 'Quest' A/L was deactivated after running for at least 6 hrs. (Per Flight Rule, when A/L CCAA is activated following a seven-day inactive period, it must be operated a minimum of 6 hrs continuously without water carryover before being shut down. The CCAA may thereafter remain inactive for 7 days with the LTL (Low Temperature Loop) set point below the USOS (U.S. Segment) dew point.)
Weekly Science Update (Expedition Seventeen -- Week 10)
3-D SPACE : After a successful session on 6/13, the second session for the FE-2 is currently planned for 6/30.
ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Measurements continue in FGB module.
ANITA (Analyzing Interferometer for Ambient Air): Continuing.
BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.
CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.
CFE (Capillary Flow Experiment): Reserve.
CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Samples returned on 1J.
CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): The CGBA Ops team is hoping to run Side 1 of Silicate Garden Hab (S/N 2) in the next couple of weeks to collect data for the upcoming school year. No crew action needed.
CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.
CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.
EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.
ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.
EPO (Educational Payload Operations): Reserve.
ETD (Eye Tracking Device): Completed.
EuTEF (European Technology Exposure Facility): The experiment platform continues to operate nominally. The instruments and its power feeder #1 have been de-activated on 6/21 because of planned power outage in ERASMUS USOC. Afterwards the EuTEF instruments have been re-activated nominally. A DHPU (Data Handling Processing Unit) software patch is currently being developed to fix the link error issues encountered with DEBIE-2 and FIPEX instruments. -- DEBIE-2: Link error still in work. However a work-around allows for regular science data acquisition using an on-board IOP (Instrument Operations Procedure); -- DOSTEL: On-going science acquisition; -- EuTEMP: Currently inactive as planned; -- EVC: still under investigation. On 0/23, a link with the HRM could successfully be established; -- EXPOSE: On-going science acquisition; -- FIPEX: It was noticed that a script run during Week 10 had to be repeated as the data file was corrupted. The file has been deleted and a new script has been started on 6/25, containing fast measure program with 3 sensors; -- MEDET: Some commanding was performed on 6/23 to increase the frequency of acquisition of the microcalorimeter. Commands to bring back to the initial value were sent on 6/24. Continuously acquiring science data; -- PLEGPAY: currently in READY mode, no science data acquisition on-going; -- TRIBOLAB: The Ball-Bearing experiment 1/2/3 which is running since 6/16 has been paused during the EUTEF power-down on 6/21 and has been resumed since. The experiment run is expected to last until ~7/18-20.
FSL (Fluid Science Laboratory): Further troubleshooting activities on FSL are planned for 7/1.
GEOFLOW: The start of GEOFLOW is pending further FSL troubleshooting and remaining commissioning activities (check of FSL optical modes).
IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.
InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.
Integrated Immune: In progress.
KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.
LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Node-1 surfaces were expected to contain little/no fungal material, but Greg's 2nd swab contained 18 ng/ml of the fungal cell wall molecule 'Glucan' (within the 1-100 ng/ml detection range of the instrument). This level is not a concern for crew health on the ISS and probably represents dead residual fungal material killed during cleaning. However, such a reading might be a scientific concern if obtained during a human exploration mission to Mars, where biological material on internal surfaces could be transferred to the outside, e.g. during EVA. This transfer could interfere with experiments designed to search for biological molecules indigenous to the Martian surface. Chamitoff's ISS results are helping scientists define requirements for spacecraft 'cleanliness', which need to be in place for these exploration missions.
MISSE (Materials ISS Experiment): Ongoing.
MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 'Pirs'.
MSG-SAME (Microgravity Science Glovebox): Complete.
NOA-2 (Nitric Oxide Analyzer): Planned.
NUTRITION w/REPOSITORY: In progress.
PADLES (Passive Dosimeter for Lifescience Experiment in Space): In progress.
PMDIS (Perceptual Motor Deficits in Space): Complete.
SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.
SAMPLE: Inc16 samples have been downloaded with 15S and handed over to science team.
SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): 'Greg, thanks for completing your Actiwatch initialization. If you find your other Actiwatch, please let us know so we can coordinate the downloads of both Actiwatches into the timeline for your next session.'
SOLAR (Solar Monitoring Observatory): Sun observation window was expected to open on 6/24, however the sun was not sufficiently visible to start measurements. On 6/25, the SOLAR platform has been reconfigured to PM (Pointing Mode). After a while, however, the HK data were lost. Some test on 6/25 and power cycles on 6/25 & 26 have not resolved the problem. The issue appears to be related either to the LAN or to an internal process of SOLAR not sending HK packets. The instruments are currently switched off. Currently under further investigation.
SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.
Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.
TRAC (Test of Reaction & Adaptation Capabilities): Planned.
WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): In progress.
CEO (Crew Earth Observations): Continuing.
CEO (Crew Earth Observations) photo targets uplinked for Greg Chamitoff for today were Mount Rainier (detailed images of Mt. Rainier, right of track, were requested. This active volcano lies immediately behind Seattle), Yellowstone National Park, Wyoming (detailed images were requested for information about roads and trails. Mapping series of overlapping images left of track was requested. Yellowstone Lake was in the center of the view and the NP), and South Amazonian Fans, Brazil (this site is almost never called out due to cloudiness (forecast was mostly clear for today). The Aripuana River and its neighbor in central Amazonia have created vast fan-shaped features ('inland deltas,' - polygon). Research is centering on old channels which appear as a series of sinuous white lines: aiming the lens left and right of track to acquire a mapping swath of images. Visual cues were a highway and the Madeira River on one side and the Trans-Amazon Highway on the other. JSC geologists are investigating inland deltas worldwide as analogs for similar features on Mars).